A 10 newton force acts on an object for 0.01 seconds. What force, acting on the object for 0.05 seconds, would produce the same impulse?
1
Expert's answer
2019-01-07T09:04:00-0500

By the definition of the impulse, we get:

"J = F\u2206t,"

here,

"J"

is the impulse,

"F"

is the force acting on the object and

"\u2206t"

is the time interval at which the force acts on the object.

Then, from this formula we can find the impulse which produced by the

"10 N"

force acting on the object for 0.01 seconds:

"J = F∆t = 10 N \cdot 0.01 s = 0.1 N \cdot s."

Then, the time at which the force acts on the object was change to 0.05 seconds. The force , acting on the object can be found from the same formula:

"F = \frac{J}{∆t} = \frac{0.1 N \cdot s}{0.05 s} = 2 N."

Answer:

"F = 2 N"

.